Two forces of F =⟨75/√2 ,−75/√2⟩ N and F =⟨150/√2 ,−150/√2⟩ N act on an object. Find the third force F3 that is needed to balance the first two forces. 


1
Expert's answer
2021-06-13T07:09:17-0400

"F'=F_1+F_2=⟨\frac{225}{\sqrt 2},~-\frac{225}{\sqrt 2}⟩,"

"F_3=-F'=⟨-\frac{225}{\sqrt 2},~\frac{225}{\sqrt 2}⟩."